3.24. The visual threshold of the normal human eye is about 7.3 x 10-11 Wcm- lor light whose wavelength is 556 nm. What is the corresponding photon flux, in photons per square centimeter per second?
Answer #1

H = 7.3×10-15W/cm^2

Wavelength = 556nm = 556×10-9m = 556×10^-7cm

Now energy E = hc/wavelength = 6.626×10^-34 × 3×10^8 / 556×10^-7

=) E = 0.035752 × 10^-19 Joule

Now photon flux = H/E = 7.3×10^-15/0.035752×10^-19

=) Flux = 2041843.813 photons per second per square centimeters.

Also flux = 2.04×10^6 photons per square cm per second .

Here I took H = 7.3×10^-15w/cm^2

In question it is not clearly visible. So in case of showing wrong answer please confirm me value of H , I'll correct .